OpenStudy (lilai3):

@waterineyes Mrs. Jackson gave her son Ronald some money to spend on clothes. Ronald was alllowed to spend up to $450. Ronald was only interested in buying jeans and shoes. The jeans cost $55 each and the shoes cost $110 each. He needs to buy at least one more pair of jeans than shoes.

OpenStudy (anonymous):

After you.. :P

OpenStudy (lilai3):

so i got... 55j+110s<=450 and s+1<=j

OpenStudy (anonymous):

First equation is correct..

OpenStudy (anonymous):

And second one too..

OpenStudy (lilai3):

oh i trust you this time too... and okay.thank you. one more problem is needed help with. please hellp

OpenStudy (anonymous):

FIrst one is same as I explained earlier that this is case of total money.. For Jeans we spent : Cost*Number of Jeans = 55j Similarly shoes : 110s And total cost exceeds no more than 450: 55j + 110s <= 450

OpenStudy (anonymous):

And there is now one condition that : Jeans pair will be atleast one more than shoes : ie Jeans pair >= 1 + shoes pair implies j >= 1 + s

OpenStudy (anonymous):

Getting??

OpenStudy (anonymous):

Or you can write equation second simply as : j - s >= 1..

OpenStudy (lilai3):

thank you can you help me with one last question. this one is sorta confusing

OpenStudy (lilai3):

i'll tag you and close this one

OpenStudy (anonymous):

Hey don't.. I am going to brush my teeth :)

OpenStudy (lilai3):

awwww fine i'll see you thank you a lot anyway

OpenStudy (anonymous):

Last one I am solving for you. go to that post and reply..